OpenStudy (anonymous):

If the range of the function y = 10x is all positive real numbers, the range of its inverse would be: Select one: a. All real numbers b. Negative Integers only c. All positive real number only d. None of the above

OpenStudy (anonymous):

i think ur function should be \(\Huge y=10^x\) right ?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok , then whats the inverse?

OpenStudy (anonymous):

y=10^x log y =x y=log x ( is the inverse function of base 10 )

OpenStudy (anonymous):

now range of log funtion of base 10 , what it is ?

OpenStudy (anonymous):

is the Ansewr A all real numbers

OpenStudy (anonymous):

yeah ^^